OpenStudy (anonymous):

Find the axis of the graph of the function. f(x) = 4x2 + 8x + 7 x = 3 x = -1 x = -2 x = 2

OpenStudy (tkhunny):

"Completing the Square" Look it up. Read about it. Practice it. Get good at it. f(x) = 4x^2 + 8x + 7 f(x) = (4x^2 + 8x) + 7 f(x) = 4(x^2 + 2x) + 7 f(x) = 4(x^2 + 2x + _______) + 7 - 4(______) 2/2 = 1 1^2 = 1 f(x) = 4(x^2 + 2x + 1) + 7 - 4(1) f(x) = 4(x+1)^2 + 7 - 4 f(x) = 4(x+1)^2 + 3 It is now relatively obvious how to answer the question - IF you know how to see it in this form.